Solve the following kinematic and free-fall problems. Show your

complete solution and box your final answer. Be sure to have a unit for your final answer and direction, when needed.





11. A phone falls from a bridge after being slapped away in the trend of ‘smack cam’ videos. The

phone falls from the bridge with an initial downward velocity of 0.4 m/s. (6pts)


a. How far would it fall when it’s been falling for 2 seconds?


b. If it fell for 50m, what would be its final velocity?


1
Expert's answer
2021-10-19T15:17:11-0400

a. The distance is


"d=v_0t+gt^2/2,\\
d=0.4·2+9.8·2^2/2=20.4\text{ m}."

b. The final velocity is


"v=\sqrt{v_i^2+2gh}=31.3\text{ m/s}."